OpenStudy (anonymous):

What is the solution of the system? Use elimination. 3x – 4y = –24 x + y = –1

OpenStudy (anonymous):

from second equation you have x =- 1 - y , plug this in the first equation and you will have your solution

OpenStudy (anonymous):

x = -y - 1 3(-y -1) - 4y = -24 Solve that to get y, then plug that into the second equation to get x.